Calculate the energy of neutrons that is seen for a Bragg angle of 45 degree for a first order reflection if the brass space 1.22A
1
Expert's answer
2019-07-08T10:26:43-0400

By Bragg-Wulf condition

"\lambda = \sqrt[2]2d"

respectively

"\epsilon = \frac{h^{2}}{4d^{2}m}= \frac{29.5\cdot10^{-28}}{6.68} = 4.4\cdot10^{-21} (J)"